# PowerGraph: A power grid benchmark dataset for graph neural networks

### v3

@@ 1 Introduction
  • A dataset that contributes to the society domain within the OGB graph dataset taxonomy [18].
@@ 3 Benchmarking the PowerGraph dataset
For each power grid dataset, both for graph and node-level tasks we utilize four baseline GNN architectures: GCNConv [27], GATConv [28], GINEConv [29], and TransformerConv [30], which are commonly used in the graph xAI community. We experiment with state-of-the-art graph transformer convolutional layers [30], forming the backbone of recent models such as GraphGPS [31], Transformer-M [32], and TokenGT [33], chosen for their ability to incorporate edge features relevant to power grids. The GNN architecture includes multiple message passing layers (MPLs), each followed by a PReLU activation function [34]. When graph-level tasks are involved, a maximum global pooling operator for obtaining graph-level embeddings from node embeddings, and one fully connected layer. We perform a grid search over the number of MPLs (1, 2, 3) and the hidden dimensionality (8, 16, 32). The Adam optimizer is used with an initial learning rate of $10^{-3}$, and each model is trained for 200 epochs, with the learning rate adjusted using a scheduler that automatically reduces it if a metric stops improving. The datasets are split into 85% train, 5% validation, and 10% test sets, with a batch size of 32 for node-level tasks and 16 for graph-level tasks. For classification models, balanced accuracy [35] is used due to class imbalance, (see Table 3). For regression models, mean squared error is used as the reference metric.
Figure 3: Node-averaged Mean Absolute Errors on the predicted physical quantities for the power flow and optimal power flow problems.
@@
Table 4 presents the best models for node-level tasks using GCN, GAT, GINe, and Transformer architectures across four power systems. Similarly, Table 5 reports the top-performing models for graph-level tasks with the same architectures and various hyperparameter combinations. In addition to the presented result in this paper and its appendices, we present the complete results, including all models and their performances, on our website: https://powergraph.ivia.ch. The website also includes Gradient Boosted Trees results as a baseline for comparison. Figure 3 illustrates the node-averaged MSE for each predicted physical quantity.
@@ Discussion
We train our models using data from current simulation methods (see A.2.1 and A.3), aiming for the highest accuracy in classification and regression to replicate those results, while the key advantage of our ML models lies in surpassing traditional solvers or simulations in computational efficiency. Concerning node-level tasks, the OPF task is inherently more complex, resulting in a higher mean squared error (MSE) for the ML model trained on it compared to the one trained on the PF task, as illustrated in Figure 3, which presents the node-averaged MSE for each predicted physical quantity. While the global MSE is consistently lower for the PF task, the PF models do not always outperform the OPF models in predicting voltage magnitude and angle, particularly for the IEEE118-bus system. In the PF problem, where generated power is known except for the slack bus, predictions for $P_{g}$ and $Q_{g}$ are more accurate. However, the models on the IEEE39 and IEEE118 grids show the worst performance across both tasks. Conversely, the PF model performs best on the UK grid, though the MSE increases in the OPF task, whereas the IEEE24 grid models show strong performance in both tasks. The fewer nodes in the IEEE24 and UK grids simplify node-level regression tasks.
@@
Notably, for node-level tasks, the GAT model generally outperforms others. Nevertheless, the GCN model, which disregards edge features, also shows strong results without significantly impacting performance. The best results are consistently achieved with a single MPL across all cases. For graph-level tasks, the GCN model consistently shows lower performance compared to other models. This highlights a clear distinction from power flow tasks, where edge features like line limits are crucial for modeling potential cascades and therefore have a significant impact on performance. In contrast, the Transformer and GINe models emerge as the best performers. The Transformer model leverages its attention mechanism to dynamically weigh the significance of neighboring nodes, capturing complex relationships and leading to superior learning outcomes. This attention-based approach is a key factor in the Transformer’s and GAT’s effectiveness across both node-level and graph-level tasks. Additionally, the GINe model demonstrates robust performance due to its strong capability in capturing and aggregating local neighborhood information, which enhances its predictive accuracy [29].
@@
While binary and multi-class classification models show good results, the regression model predicting the exact demand not served does not. The R2 [36] score, also known as the coefficient of determination, reaches a maximum of 0.43 for the best regression model on the IEEE24 dataset, but remains below 0.26 for other power systems. Ideally, an R2 score closer to 1 is desired for a better model fit. This highlights the need for future research to focus on improving GNN performance on regression tasks both at the graph and node-level. We observe the limitations of GNNs, particularly in node regression tasks, where simpler models like Gradient Boosted Trees (GBT) outperform GNNs (see https://powergraph.ivia.ch). However, this trend does not extend to graph-level tasks, where GNNs consistently prove to be the superior method. This highlights the need for further development of GNN models tailored specifically for node regression tasks, which are critical in power systems analysis. At the same time, even minor improvements in classification tasks should not be underestimated, as they can significantly impact decision-making in critical infrastructures, as power grids.
@@
IEEE24 Powerflow gat 1h 32n 1.12e-02±3.34e-04
Optimal powerflow gat 1h 8n 1.71e-02±3.44e-04
IEEE39 Powerflow gcn 1h 16n 1.94e-02±2.19e-04
Optimal powerflow gat 1h 16n 2.14e-02±1.55e-03
IEEE118 Powerflow gat 1h 32n 1.12e-02±3.34e-04
Optimal powerflow gat 1h 8n 1.71e-02±3.44e-04
UK Powerflow transformer 1h 32n 1.01e-02±2.47e-03
Optimal powerflow gat 1h 32n 1.55e-02±3.27e-03
@@ 6 Conclusions
PowerGraph addresses critical challenges encountered in utilizing Graph Neural Networks (GNNs) for power systems analyses. PowerGraph effectively fills a significant gap in the GNN domain for power engineering applications by offering a dataset tailored for node and graph-level tasks and model explainability. Through rigorous benchmarking against a range of GNN and explainability models, PowerGraph exhibits high performance in graph classification, albeit indicating a need for further refinement in regression models. Regression models remain essential in power systems and can be solved by GNNs for tasks like power flow and system security analysis. However, GNN still encounters challenges that demand further research and development. Additionally, the complexity of architecture varies considerably between node-level and graph-level tasks. For instance, power flow analysis at the node level requires fewer message-passing layers compared to cascading failure analysis at the graph level. In power flow analysis, we conclude that nearest neighbor information is crucial for accurate output prediction. However, this is not the case for predicting cascading failures, which requires information from more distant nodes for accurate predictions. PowerGraph is the first real-world dataset to offer ground-truth explanations for graph-level tasks in explainable AI, enabling thorough evaluation of various explainability methods. Despite this advancement, the performance of these methods remains suboptimal, highlighting the lack of a universally superior approach. Given the crucial role of explainability for power grid operators, this underscores the ongoing need for dedicated research and development in this field. Looking forward, we plan to enhance PowerGraph by adding a temporal graph dataset to facilitate in-depth analysis of the stages of cascading failures. Although this information is already present in the dataset, please refer to A.5, it was beyond the scope of this work. We will also benchmark methods on larger synthetic power systems, such as the grid in [49], to assess whether deeper GNN architectures are necessary, as noted in [50]. A raw dataset is available at https://powergraph.ivia.ch.

### v4

@@ 3 Benchmarking the PowerGraph dataset
For each power grid dataset, both for graph and node-level tasks we utilize four baseline GNN architectures: GCNConv [27], GATConv [28], GINEConv [29], and TransformerConv [30], which are commonly used in the graph xAI community. We experiment with state-of-the-art graph transformer convolutional layers [30], forming the backbone of recent models such as GraphGPS [31], Transformer-M [32], and TokenGT [33], chosen for their ability to incorporate edge features relevant to power grids. The GNN architecture includes multiple message passing layers (MPLs), each followed by a PReLU activation function [34]. When graph-level tasks are involved, a maximum global pooling operator for obtaining graph-level embeddings from node embeddings, and one fully connected layer. We perform a grid search over the number of MPLs (1, 2, 3) and the hidden dimensionality (8, 16, 32). The Adam optimizer is used with an initial learning rate of $10^{-3}$, and each model is trained for 50 epochs, with the learning rate adjusted using a scheduler that automatically reduces it if the reference metric stops improving. The datasets are split into 85% train, 5% validation, and 10% test sets, with a batch size of 32 for node-level tasks and 16 for graph-level tasks. For classification models, balanced accuracy [35] is used due to class imbalance, (see Table 3). For regression models, mean squared error is used as the reference metric.
Figure 3: Node-averaged Mean Absolute Errors on the predicted physical quantities for the power flow and optimal power flow problems on the best performing models reported in Table 4.
@@
Table 4 presents the best models for node-level tasks using GCN, GAT, GINe, and Transformer architectures across the four power systems. Similarly, Table 5 reports the top-performing models for graph-level tasks. In addition to the presented result in this paper and its appendices, we present the complete results, including all models and their performances, on our website: https://powergraph.ivia.ch. The website also includes Gradient Boosted Trees results as a baseline for comparison. Figure 3 illustrates the node-averaged MSE for each predicted physical quantity for the power flow tasks for the best performing models reported in Table 4.
@@ Discussion
We train our models using data from current simulation methods (see A.2.1 and A.3), aiming for the highest accuracy in classification and regression to replicate those results, while the key advantage of our ML models lies in surpassing traditional solvers or simulations in computational efficiency. For node-level tasks, the OPF task presents a higher level of complexity, leading to a higher mean squared error (MSE) for the machine learning model trained on it compared to the model trained on the PF task, as shown in Figure 1 and Table 4. Although model performance varies slightly across different grids, we observe that GNNs are effective in predicting optimal power flow and power flow solutions across grids with varying topologies, achieving low prediction error for key physical quantities regardless of grid size.
@@
For both graph-level and node-level tasks, the Transformer model consistently outperforms other approaches, achieving optimal results, particularly with three MPL layers in most cases. In contrast, the GCN model repeatedly demonstrates lower performance compared to other models, emphasizing its limitations for power flow tasks where edge features are critical for analyzing power flow and modeling potential cascades. This highlights the advantages of the Transformer and GINe models, which emerge as the top performers. The Transformer’s attention mechanism allows it to dynamically assign importance to neighboring nodes, capturing intricate relationships that enhance its predictive accuracy. This attention-based approach significantly contributes to the effectiveness of both the Transformer and GAT models across node-level and graph-level tasks. Meanwhile, the GINe model also shows strong performance due to its powerful ability to capture the graph topology of power grids, further boosting its predictive accuracy [29].
@@
While binary and multi-class classification models show good results, the regression model predicting the exact demand not served does not. For graph-level regression tasks, the R2 [36] score, also known as the coefficient of determination, peaks at 0.43 for the best model on the IEEE24 dataset but falls below 0.26 across other power systems. Ideally, an R2 score closer to 1 is desired for a better model fit, underscoring the need for further research to enhance GNN performance on regression tasks. on improving GNN performance on regression tasks. Additionally, GNNs show limitations on node-level regression, where simpler models like Gradient Boosted Trees (GBT) sometimes outperform them on certain datasets (see https://powergraph.ivia.ch). However, this trend does not extend to graph-level tasks, where GNNs consistently prove to be the superior method. This calls for the development of GNN architectures specifically optimized for regression tasks, crucial for power systems analysis. At the same time, even minor improvements in classification tasks should not be underestimated, as they can significantly impact decision-making in critical infrastructures such as power grids.
@@
IEEE24 Powerflow transformer 3h 32n 4.35e-05±7.46e-06
Optimal powerflow transformer 3h 32n 8.99e-05±3.43e-06
IEEE39 Powerflow transformer 3h 32n 4.35e-05±1.27e-05
Optimal powerflow transformer 3h 32n 7.38e-05±7.87e-06
IEEE118 Powerflow transformer 3h 32n 2.39e-05±3.74e-06
Optimal powerflow transformer 3h 32n 5.74e-05±4.24e-06
UK Powerflow transformer 3h 32n 2.02e-05±8.68e-06
Optimal powerflow transformer 3h 32n 3.75e-05±1.39e-05
@@ 6 Conclusions
PowerGraph addresses critical challenges encountered in utilizing Graph Neural Networks (GNNs) for power systems analyses. PowerGraph effectively fills a significant gap in the GNN domain for power engineering applications by offering a dataset tailored for node and graph-level tasks and model explainability. Through rigorous benchmarking against a range of GNN and explainability models, PowerGraph exhibits high performance in graph classification, albeit indicating a need for further refinement in regression models. Regression models remain essential in power systems and can be solved by GNNs for tasks like power flow and system security analysis. However, GNN still encounters challenges that demand further research and development. Across power grid analysis tasks, we find that an architecture with three message-passing layers yields the most accurate predictions for both power flow and cascading failure analysis. This result suggests that incorporating information from up to three-hop neighbors is essential for precise modeling, effectively balancing local and global network information required for accurate predictions in both graph- and node-level tasks. Furthermore, PowerGraph is the first real-world dataset to offer ground-truth explanations for graph-level tasks in explainable AI, enabling thorough evaluation of various explainability methods. Despite this advancement, the performance of these methods remains suboptimal, highlighting the lack of a universally superior approach. Given the crucial role of explainability for power grid operators, this underscores the ongoing need for dedicated research and development in this field. Looking forward, we plan to enhance PowerGraph by adding a temporal graph dataset to facilitate in-depth analysis of the stages of cascading failures. Although this information is already present in the dataset, please refer to A.5, it was beyond the scope of this work. We also benchmark methods on larger synthetic power systems, such as the grid in [49], to assess whether deeper GNN architectures are necessary, as noted in [50]. A raw dataset and the benchmarking results are available at https://powergraph.ivia.ch.